Graph the line passing through the given point with the given slop.
(-2,2), -1/3

Found 2 solutions by drk, MRperkins:
Answer by drk(1908)

M = -1/3
Y = mx + b
By substitution, we get
2 = (-2)*(-1/3) + b
4/3 = b
Y = (-1/3)x + (4/3)

Answer by MRperkins(300)

find the point (-2,2) and then go down one and right 3. Connect the dots.
remember that the coordinate (-2,2) is in the form (x,y) so find where x is -2 and y is 2 and the point where they interesect is the point (-2,2).